Could someone please help me solve this equation? One canned juice is 15% orange juice, another is 10% orange juice.How many liters of each should be mixed together in order to get 5 liters that is 11% orange juice?
Found 2 solutions by stanbon, rwm:
Answer by stanbon(75887)

One canned juice is 15% orange juice, another is 10% orange juice.How many liters of each should be mixed together in order to get 5 liters that is 11% orange juice?
--------
Equation:
0.15x + 0.10(5-x) = 0.11*5
---
Multiply thru by 100:
15x + 10*5 - 10x = 11*5
---
5x = 5
x = 1 liter (amt. 15% oj needed)
----
5-x = 4 liters (amt of 10% oj needed)
===========================================
cheers,
Stan H.
=======================

Answer by rwm(914)

.15x+.10*y=5*.11
x+y=5
x=1
y=4